# Statistical reduced order modelling for the parametric Helmholtz equation

### v1

@@ Abstract
Predictive modeling involving simulation and sensor data at the same time, is a growing challenge in computational science. Even with large-scale finite element models, a mismatch to the sensor data often remains, which can be attributed to different sources of uncertainty. For such a scenario, the statistical finite element method (statFEM) can be used to condition a simulated field on given sensor data. This yields a posterior solution which resembles the data much better and additionally provides consistent estimates of uncertainty, including model misspecification. For frequency or parameter dependent problems, occurring, e.g. in acoustics or electromagnetism, solving the full order model at the frequency grid and conditioning it on data quickly results in a prohibitive computational cost. In this case, the introduction of a surrogate in form of a reduced order model yields much smaller systems of equations. In this paper, we propose a reduced order statFEM framework relying on Krylov-based moment matching. We introduce a data model which explicitly includes the bias induced by the reduced approximation, which is estimated by an inexpensive error indicator. The results of the new statistical reduced order method are compared to the standard statFEM procedure applied to a ROM prior, i.e. without explicitly accounting for the reduced order bias. The proposed method yields better accuracy and faster convergence throughout a given frequency range for different numerical examples.
@@ 1 Introduction
In engineering and physics, but also in many other disciplines such as chemistry and finance, partial differential equations (PDEs) [1] serve as a very fundamental mathematical description of a given problem. Generally, most PDEs are hard or impossible to solve analytically. To find an approximation of the solution, numerical methods such as the finite element method (FEM) [2] are used. Both the model assumptions, i.e. choice of the differential operator, parameters and boundary conditions, which lead to the PDE and the approximative nature of the numerical method contribute to a discrepancy between the solution of the numerical model and the physical reality. Collecting data with sensors may yield relatively accurate, albeit noisy, observations at discrete locations in the computational domain. The numerical solution, on the other hand, yields a continuous solution at every point in the domain. Combining the numerical solution with the sparse observation data to estimate the state of a system is referred to as data assimilation [3]. One data assimilation technique is the recently proposed statistical FEM (statFEM) [4]. In statFEM, a statistical model for measured sensor data is introduced in which the data is additively composed of the numerical FE solution, a model error and measurement noise. All these terms are modelled as Gaussian Processes (GP) [5], which enables data assimilation through Bayesian principles. The stochastic solution of the numerical problem is considered as a so-called Bayesian prior. This prior can be conditioned on measurement data to lead to an improved solution, called the posterior. Accounting for uncertainty in the FE model, a Gaussian prior for the state is obtained by first-order-second moment analysis. Despite acknowledging the uncertainty, the model may still be misspecified, in particular for complex physical processes. Hence, a model error GP is introduced, which is assumed to be mean-free, i.e. a possible discrepancy between model and reality will be considered through a higher variance of the model error. Adopting an empirical Bayes approach, the hyperparameters of the covariance function are learned from the data first, before the solution GP is conditioned on the data, where measurement error is also accounted for.
@@
Some engineering problems require solving PDEs depending on deterministic parameters in addition to the uncertain model inputs. This concerns frequency domain problems, i.e. the Helmholtz equation in acoustics [6], Maxwell’s equations in electromagnetism [7] or the equations of motion in structural dynamics [8]. Concerning the latter, possible use cases can be found e.g. in the domain of structural health monitoring, where the state of e.g. a large scale structure [9, 10] or smaller scale aerospace turbine components is constantly observed by means of sparsely located sensors and a structural dynamics model which has to be updated for the full frequency range once new measurements come in [11]. The resulting FEM models can become very complex and expensive to solve even for a single frequency. Hence, computing a parameter sweep, i.e. solving the system at many frequencies within a desired interval, quickly becomes prohibitively expensive. Through reduced order models (ROM), the solution of e.g. a FEM system of equations with a large number of degrees of freedom (DOF) can be approximated by solving a significantly smaller ROM system of equations. The most prominent techniques for model order reduction are projection based ROM methods [12]. Examples of those are proper orthogonal decomposition (POD) or Krylov subspace methods. As, in general, neither of these ROM methods are a perfectly accurate surrogate for the FEM, also the ROM approximation error has to be taken into account, see for instance [13]. In contrast to the model to reality error, this ROM approximation error has a well-known structure, governed by the error, or residual, equation.
@@
Here, an adjoint based error estimator [14, 15] is derived which doesn’t involve solving the full order system matrix. This additional error is then included into the data model. In principle, the ROM error could also be estimated as part of the model error. Here, this will be referred to as the classical approach. However, as it is possible to estimate the ROM error without solving the full system, it seems sensible to separate it from the general model error and hence have a better prior for the total error. In this paper, a statistical reduced order modeling approach, referred to as statROM, is introduced. It is shown that this approach leads to a better approximation of the data generating process and faster convergence of the error. The method is applied to a FEM approximation of the Helmholtz equation in 1D with few DOFs to study the convergence properties in much detail. Data are generated from a fine FEM discretisation. The reduced order model is constructed with a moment matching Krylov subspace method. Next, a 2D Helmholtz scattering problem with significantly more DOFs is evaluated to study how the method scales.
@@ 2 Model problem
We consider the Helmholtz equation as an example of a frequency dependent PDE, which serves as a model problem, e.g. for wave-based non-destructive testing applications. More specifically, the model PDE is given as
$\Delta u(\bm{x})+k^{2}\kappa(\bm{x})u(\bm{x})$ $=-f(\bm{x}),\quad$ in $\text{in }D,$ (1)
$u(\bm{x})$ $=d(\bm{x}),\quad$ on $\text{on }\Gamma_{\mathrm{D}},$ (2)
$\frac{\partial u(\bm{x})}{\partial n}$ $=g(\bm{x}),\quad$ on $\text{on }\Gamma_{\mathrm{N}},$ (3)
$\frac{\partial u(\bm{x})}{\partial n}-\mathrm{i}k\beta u(\bm{x})$ $=0,\quad$ on $\text{on }\Gamma_{\mathrm{I}},$ (4)
on a bounded domain $D\subset\mathbb{R}^{d}$ with $d=1,2,3$, where $\partial D$ is a Lipschitz boundary for $d\geq 2$ that can be approximated exactly by polygons or polyhedra. A sound soft boundary condition can be realized with $d(\bm{x})=0$. The Neumann boundary can be separated into a homogeneous, i.e. sound hard reflecting part $\Gamma_{\mathrm{N},0}$ and an inhomogeneous part $\Gamma_{\mathrm{N},g}$, where $g\neq 0$. Damping can be considered by introducing the Robin boundary condition on $\Gamma_{\mathrm{I}}$, where $\beta\in\mathbb{R}$. Note that $\Gamma_{\mathrm{D}}$ and/or $\Gamma_{\mathrm{I}}$ may be empty. Furthermore, there holds $k\in\mathbb{R}=\frac{\omega}{c}$, for the wave number, with $\omega$ denoting the angular frequency and $c$ the speed of sound. Several quantities in the model problem are assumed to be unknown and hence, exhibit uncertainty. These are the material parameter $\kappa$, the inhomogeneity in the Neumann boundary condition $g$ and also the forcing $f$.
@@
[…]
@@
[…]
@@
with $\bm{M}$ the mass matrix, $\bm{D}$ the damping matrix and $\bm{S}$ the stiffness matrix. In (13), $\bm{A}\in\mathbb{C}^{N\times N}$ denotes the system matrix, $\bm{u}\in\mathbb{C}^{N}$ the nodal coefficients and $\bm{f}\in\mathbb{R}^{N}$ the source vector.
At this stage, we also introduce random influences, which model unknown material parameters and forcings. More precisely, $\kappa,f,g$ will be modelled as a random fields on a common probability space $(\Theta,\Sigma,\mathcal{P})$ and we obtain
$\bm{A}(\omega,\theta)\bm{u}(\omega,\theta)=\bm{f}(\theta),\ \forall\omega\in[\omega_{\text{min}},\omega_{\text{max}}]\subset\mathbb{R}_{+}$ (14)
@@ 3 Reduced Order Modelling
In many situations in practice, the size of the discrete system $N$ is very large, which makes it impossible to solve problem (13) for many different values of the angular frequency $\omega$ or to efficiently perform an uncertainty analysis. Therefore, from the FEM system of equations, a Reduced Order Model (ROM) is typically derived, which can be used for multi-query tasks. As the model problem is defined in frequency domain, an uncertainty analysis is especially expensive, because it needs to be repeated for each point in the frequency domain. With a suitable ROM that describes the underlying FE-problem accurately, a tremendous reduction of numerical cost can be achieved, which in turn enables time-critical applications where, for instance, the model has to be conditioned on new incoming sensor data very frequently.
@@
FEM $:\quad\bm{A}(\omega)\bm{u}(\omega)=\bm{f},$ (15)
ROM $:\quad\bm{A}_{r}(\omega)\bm{u}_{r}(\omega)=\bm{f}_{r},$ (16)
where $\bm{A}_{r}(\omega)\in\mathbb{C}^{r\times r},\bm{u}_{r}(\omega),\bm{f}_{r}\in\mathbb{C}^{r}$ and $r\ll N$. The FE solution and the ROM solution vector are connected by a projection matrix $\bm{V}\in\mathbb{C}^{N\times r}$, as
$\bm{u}(\omega)\approx\bm{V}\bm{u}_{r}(\omega)=\tilde{\bm{u}}(\omega),$ (17)
as well as $\bm{A}_{r}=\bm{V}^{*}\bm{A}\bm{V}$ and $\bm{f}_{r}=\bm{V}^{*}\bm{f}$. For the projection matrix there holds $\bm{V}^{*}\bm{V}=\bm{I}$, hence it is orthogonal.
@@
A class of projection based ROMs which is especially well suited for dynamical systems [20] are the moment matching methods. We adopt the adaptive order rational Arnoldi-type method (AORA) as proposed in [21] and employ in particular the formulation for second order systems as put forth in [20]. Moment matching is based on rational interpolation [22] and the resulting ROM satisfies the Hermite interpolation conditions up to some order [12]. To generate a projection matrix $\bm{V}$ which has the moment matching property, the system components of the dynamical 2nd order system (12) are first used to derive a generalized Krylov subspace
$\mathcal{K}(\bm{M},\bm{D},\bm{S},\bm{f})=G_{k}(-\bm{A}(\bar{\omega})^{-1}\bm{A}^{\prime}(\bar{\omega}),-\bm{A}(\bar{\omega})^{-1}i\bm{M},\bm{A}(\bar{\omega})^{-1}\bm{f})$ (18)
@@
$G_{k}(\bm{L},\bm{B},\bm{s})=\mathrm{span}\{\bm{s}_{0},\bm{s}_{1},\dots,\bm{s}_{m-1}\},\;\bm{s}_{0}=\bm{s},\bm{s}_{1}=\bm{L}\bm{s}_{0},\,\\ \text{and}\,\bm{s}_{l}=\bm{L}\bm{s}_{l-1}+\bm{B}\bm{s}_{l-2}\,\text{for}\,l\geq 2$ (19)
in the case of only a single expansion frequency $\bar{\omega}$ and the number of matched moments $m$. Here, $\bm{A}^{\prime}(\bar{\omega})=-2\bar{\omega}\bm{M}+i\beta/c\bm{D}$. Then, one possibility to build the actual projection matrix is using a second-order rational Arnoldi-type method. This guarantees orthonormality in the columns of $\bm{V}$. In principle, Arnoldi-type methods work by transforming a matrix $\bm{A}$ into the upper Hessenberg form $\bm{H}$ by
$\bm{A}=\bm{V}\bm{H}\bm{V}^{*}\;.$ (20)
This happens iteratively and can be stopped after a certain number of iterations in order to obtain a reduced order approximation of $\bm{A}$. With the projection matrix from the rational Arnoldi method, there holds for the reduced order system
$(-\omega^{2}\bm{M}_{r}+ik\beta\bm{D}_{r}+\bm{S}_{r})\bm{u}_{r}(\omega)=\bm{f}_{r},$ (21)

### v2

@@ Abstract
Predictive modelling combining both numerical simulations and real-world measurement data, obtained from sensors, is gaining importance in computational science and engineering. Even with large-scale finite element models, a mismatch to the sensor data often remains, which can be attributed to different sources of uncertainty. For such a scenario, the statistical finite element method (statFEM) can be used to condition a simulated field on given sensor data. This yields a posterior solution which resembles the data much better and additionally provides consistent estimates of uncertainty, including model misspecification. For frequency or parameter dependent problems, occurring, e.g. in acoustics or electromagnetism, solving the full order model across the frequency range of interest and conditioning it on data quickly results in a prohibitive computational cost. In this case, the introduction of a surrogate in the form of a reduced order model (ROM) yields much smaller systems of equations. In this paper, we propose a reduced order statFEM framework relying on Krylov-based moment matching. We introduce a data model which explicitly includes the bias induced by the reduced approximation, which is estimated by an error indicator. The results of the new statistical reduced order method are compared to the standard statFEM procedure applied to a ROM prior without explicitly accounting for reduced order bias. The proposed method achieves better accuracy and faster convergence throughout a given frequency range for a variety of numerical examples and artificial (i.e. simulated) sensor data.
@@ 1 Introduction
In engineering and physics, but also in many other disciplines such as chemistry and finance, partial differential equations (PDEs) [1] serve as a very fundamental mathematical description of a given problem. Generally, most PDEs are hard or impossible to solve analytically. To find an approximation of the solution, numerical methods such as the finite element method (FEM) [2] are used. Both the model assumptions, i.e. choice of the differential operator, parameters and boundary conditions, which lead to the PDE and the approximative nature of the numerical method contribute to the difference between the solution of the numerical model and physical reality. Collecting data with sensors may yield relatively accurate, albeit noisy, observations at discrete locations in the computational domain. The numerical solution, on the other hand, yields a continuous solution at every point in the domain. Combining the numerical solution with the sparse observation data to estimate the state of a system is referred to as data assimilation [3].
One data assimilation technique is the recently proposed statistical FEM (statFEM) [4]. In statFEM, a statistical model for measured sensor data is introduced in which the data are additively composed of the numerical FE solution, a model discrepancy and measurement noise. All these terms are modeled as Gaussian Processes (GP) [5], which enables data assimilation through Bayesian principles. By propagating uncertainty through the FE model, a Gaussian prior for the state is obtained. In the original statFEM paper, a first-order perturbation approach [6] is employed, but Monte Carlo type methods are also well-suited. Once the prior is at hand, conditioning on measurement data leads to the posterior. Despite acknowledging the uncertainty, the model may still be misspecified, in particular for complex physical processes. Hence, a model error GP is introduced, which is assumed to be mean-free, i.e. a possible discrepancy between model and reality is reflected in the variance of the model error. Adopting an empirical Bayes approach, the hyperparameters of the covariance function are learned from the data first, before the solution GP is conditioned on the data, where measurement error is also accounted for.
@@
Some engineering problems require solving PDEs depending on deterministic parameters in addition to the uncertain model inputs. This concerns frequency domain problems, i.e. the Helmholtz equation in acoustics [7], Maxwell’s equations in electromagnetism [8] or the equations of motion in structural dynamics [9]. Concerning the latter, possible use cases can be found e.g. in the domain of structural health monitoring, where the state of a large scale structure [10, 11] or smaller scale aerospace turbine components is constantly observed by means of sparsely located sensors and a structural dynamics model has to be updated for the full frequency range once new measurements come in [12]. The resulting FE models can become very complex and expensive to solve even for a single frequency. Hence, computing a parameter sweep, i.e. solving the system at many frequencies within a desired interval, quickly becomes prohibitively expensive. Through reduced order models (ROM), the solution of e.g. a FEM system of equations with a large number of degrees of freedom (DOF) can be approximated by solving a significantly smaller ROM. The most prominent techniques for model order reduction are projection based methods [13]. Examples of those are proper orthogonal decomposition (POD) or Krylov subspace methods. As, in general, neither of these ROM methods are a perfectly accurate surrogate for the FEM, also the ROM approximation error has to be taken into account, see for instance [14]. In contrast to the model-to-reality error, this ROM approximation error has a well-known structure, governed by the error, or residual, equation.
@@
In this paper, a statistical reduced order modelling approach, referred to as statROM, is introduced. We show that this approach leads to a better approximation of the ground truth and faster convergence of the error. We obtain the prior by quasi Monte Carlo analysis [15] based on a frequency-dependent ROM and assimilate sparse sensor data with Bayesian updating. The ROM is constructed with a moment matching Krylov subspace method. Of particular importance in this work is the ROM error and an adjoint based error estimator [16, 17] is derived which does not involve solving the full order system of equations. This additional error is then included into the data model. In principle, the ROM error could also be estimated as part of the model error. Here, this will be referred to as the standard approach. However, as it is possible to estimate the ROM error without solving the full system, it seems sensible to separate it from the general model error and hence have a better prior for the total error.
Figure 1: Workflow diagram of the proposed statROM method. In an offline phase, the reduced order projection basis is constructed for the primal and adjoint problem and each sample of the input parameters. In an online phase the FE matrices are reduced and the system is solved in the reduced space at reduced cost compared to the full order model. The adjoint error indicator contributes to the quality of the subsequent data assimilation. Fast parameter sweeps over the frequency $\omega$ can be carried out in the online phase based on the sampled ROMs.
The method is applied to a FEM approximation of the Helmholtz equation in 1D with few DOFs to study the convergence properties in much detail. Data are generated from a fine FE discretisation. Next, a 2D Helmholtz scattering problem with significantly more DOFs is evaluated to study how the method scales.
In Figure 1, the workflow of the proposed method is sketched. A reduced order model that can be rapidly solved, enables an online phase with fast frequency sweeps. Our main contributions can be summarized as follows:
  • Extension of the statFEM method to handle parameter dependent problems with ROMs
  • Introduction of a ROM error estimator to improve the posterior accuracy
  • Application of statFEM and the proposed statROM method to the Helmholtz equation
@@ 2 Model problem
We consider the Helmholtz equation as an example of a frequency dependent PDE, which serves as a model problem, e.g. for wave-based non-destructive testing applications. In [18], an overview over such methods is provided. Specific examples in the context of structural health monitoring using high fidelity FE models can be found in e.g. [19, 20, 21]. More specifically in the case treated here, the model PDE is given as
$\Delta u(\bm{x})+k^{2}\kappa(\bm{x})u(\bm{x})$ $=-f(\bm{x}),\quad$ $\text{in }D,$ (1)
$u(\bm{x})$ $=0,\quad$ $\text{on }\Gamma_{\mathrm{D}},$ (2)
$\frac{\partial u(\bm{x})}{\partial n}$ $=g(\bm{x}),\quad$ $\text{on }\Gamma_{\mathrm{N}},$ (3)
$\frac{\partial u(\bm{x})}{\partial n}-\mathrm{i}k\beta u(\bm{x})$ $=0,\quad$ $\text{on }\Gamma_{\mathrm{I}},$ (4)
on a bounded domain $D\subset\mathbb{R}^{d}$ with $d=1,2,3$, where $\partial D$ is a Lipschitz boundary for $d\geq 2$ that can be well-approximated by polygons or polyhedra. Moreover, $\bm{n}$ denotes the outward normal to $D$. The PDE is solved for the sound pressure field $u(\bm{x})$. A sound soft boundary condition can be realized with the homogeneous Dirichlet boundary condition. The Neumann boundary can be separated into a homogeneous, i.e. sound hard reflecting part $\Gamma_{\mathrm{N},0}$ and an inhomogeneous part $\Gamma_{\mathrm{N},g}$, where $g\neq 0$. Damping can be considered by introducing the Robin boundary condition on $\Gamma_{\mathrm{I}}$, with the specific normalized acoustic admittance $\beta\in\mathbb{R}$. Note that $\Gamma_{\mathrm{D}}$ and/or $\Gamma_{\mathrm{I}}$ may be empty. Furthermore the wave number is defined as $\mathbb{R}\ni k=\frac{\omega}{c}$, with $\omega$ denoting the angular frequency and $c$ the speed of sound and $i$ is the imaginary unit. Several quantities in the model problem are assumed to be unknown and hence, exhibit uncertainty. These are the material parameter $\kappa$, the inhomogeneity in the Neumann boundary condition $g$ and also the forcing $f$.
@@
$V=H^{1}(D)=\{v\in L^{2}(D)\ |\ \|v\|_{V}^{2}=\int_{D}\nabla v\cdot\nabla v^{*}\mathrm{d}x+\int_{D}vv^{*}\mathrm{d}x<\infty\},$ (5)
@@
$s(u,v)$ $=\int_{D}\nabla u\cdot\nabla v^{*}\mathrm{d}x,$ $\text{(stiffness)},$ (8)
$m(u,v)$ $=\int_{D}\kappa uv^{*}\mathrm{d}x,$ $\text{(mass)},$ (9)
$d(u,v)$ $=\int_{\Gamma_{\mathrm{I}}}uv^{*}\mathrm{d}x,$ $\text{(damping)},$ (10)
$F(v)$ $=\int_{\Gamma_{\mathrm{N},g}}gv^{*}\mathrm{d}x+\int_{D}fv^{*}\mathrm{d}x$ $\text{(forcing)},$ (11)
@@
with $\bm{M}$ the mass matrix, $\bm{D}$ the damping matrix and $\bm{S}$ the stiffness matrix. In (13), $\bm{A}\in\mathbb{C}^{N\times N}$ denotes the system matrix, $\bm{u}\in\mathbb{C}^{N}$ the nodal coefficients and $\bm{F}\in\mathbb{C}^{N}$ the source vector.
At this stage, we also introduce random influences, which model unknown material parameters and forcings. More precisely, $\kappa,f,g$ will be modeled as random fields on a common probability space $(\Theta,\Sigma,\mathcal{P})$ with a sample space $\Theta$, a sigma algebra $\Sigma$ and probability measure $\mathcal{P}$. We obtain
$\bm{A}(\omega,\theta)\bm{u}(\omega,\theta)=\bm{F}(\omega,\theta),\ \forall\omega\in[\omega_{\text{min}},\omega_{\text{max}}]\subset\mathbb{R}_{+}$ (14)
@@ 3 Reduced Order Modelling
In many situations in practice, the size of the discrete system $N$ is very large, which makes it impossible to solve problem (13) for many different values of the angular frequency $\omega$ or to efficiently perform an uncertainty analysis. Therefore, from the FEM system of equations, a ROM is typically derived, which can be used for multi-query tasks. As the model problem is defined in frequency domain, an uncertainty analysis is especially expensive, because it needs to be repeated for each point in the frequency domain. With a suitable ROM that describes the underlying FE-problem accurately, a reduction of numerical cost can be achieved, which in turn enables time-critical applications where, for instance, the model has to be conditioned on new incoming sensor data very frequently.
@@
FEM $:\quad\bm{A}(\omega)\bm{u}(\omega)=\bm{F}(\omega),$ (15)
ROM $:\quad\bm{A}_{r}(\omega)\bm{u}_{r}(\omega)=\bm{F}_{r}(\omega),$ (16)
where $\bm{A}_{r}(\omega)\in\mathbb{C}^{r\times r},\bm{u}_{r}(\omega),\bm{F}_{r}(\omega)\in\mathbb{C}^{r}$ and $r\ll N$. The FE solution and the ROM solution vector are connected by a projection matrix $\bm{V}\in\mathbb{C}^{N\times r}$, as
$\bm{u}(\omega)\approx\bm{V}\bm{u}_{r}(\omega)$ (17)
as well as $\bm{A}_{r}=\bm{V}^{*}\bm{A}\bm{V}$ and $\bm{F}_{r}=\bm{V}^{*}\bm{F}$. Without loss of generality, the projection matrix satisfies $\bm{V}^{*}\bm{V}=\bm{I}$, hence it is orthogonal.
@@
A class of projection based ROMs which is well suited for dynamical systems [26] consists of moment matching methods. Moment matching is based on rational interpolation [27] and the resulting ROM satisfies the Hermite interpolation conditions up to some order [13]. The interpolatory approach can be considered in a structure preserving way within a much more general setting, including the second order dynamical system case as shown in [28]. In particular, since (13) is a second order dynamical system, a structure-preserving second order moment matching method analogous to the SOAR procedure [29] will be used. To do so, we will use second order generalized Krylov subspaces [30, 29] applied to (13). In the case of a single expansion frequency $\bar{\omega}$ we have that
$\mathcal{K}(\bm{M},\bm{D},\bm{S},\bm{F})=G_{m}(-\bm{A}(\bar{\omega})^{-1}\bm{A}^{\prime}(\bar{\omega}),-\bm{A}(\bar{\omega})^{-1}i\bm{M},\bm{A}(\bar{\omega})^{-1}\bm{F}(\bar{\omega}))$ (18)
@@
$G_{m}(\bm{L},\bm{B},\bm{r})=\mathrm{span}\{\bm{r}_{0},\bm{r}_{1},\dots,\bm{r}_{m-1}\},\;\bm{r}_{0}=\bm{r},\bm{r}_{1}=\bm{L}\bm{r}_{0},\,\\ \text{and}\,\bm{r}_{l}=\bm{L}\bm{r}_{l-1}+\bm{B}\bm{r}_{l-2}\,\text{for}\,l\geq 2,$ (19)
where we define the coefficient matrices $\bm{L}=-\bm{A}(\bar{\omega})^{-1}\bm{A}^{\prime}(\bar{\omega})$ and $\bm{B}=-\bm{A}(\bar{\omega})^{-1}i\bm{M}$. $\bm{r}_{0},\bm{r}_{1},\dots,\bm{r}_{m-1}$ is a second-order Krylov sequence and $m$ is the number of matched moments. For the derivative of $\bm{A}(\bar{\omega})$ with respect to the expansion frequency $\bar{\omega}$ we find $\bm{A}^{\prime}(\bar{\omega})=-2\bar{\omega}\bm{M}+i\beta/c\bm{D}$. To build the actual projection matrix $\bm{V}$ to subspace (18), we adopt the adaptive order rational Arnoldi-type method (AORA) [31] as proposed in [32] to the second order case analogous to the SOAR procedure. In this way we obtain an adaptive order second-order rational Arnoldi-type method. This ensures orthonormality in the columns of $\bm{V}$ while at the same time several expansion frequencies can be used. A central component of the method is a modified Gram-Schmidt orthogonalization. Repeating the orthogonalization once can improve the numerical stability of the method, as demonstrated in [33]. With the projection matrix $\bm{V}$ from the second order rational Arnoldi method, we obtain the following for the reduced order system:
$(-k^{2}\bm{M}_{r}-ik\beta\bm{D}_{r}+\bm{S}_{r})\bm{u}_{r}(\omega)=\bm{F}_{r}(\bar{\omega}),$ (20)